From a chemical reaction: NH3 (g) + CuO (s) ------------- N2 (g) + Cu (s) + H2O (g) Determine the limiting reagent if Ammonium has 18.1 grams and Copper Oxide has 90.4 grams>
1
Expert's answer
2020-11-23T06:49:59-0500

2NH3+3CuO=N2+3Cu+3H2O

18.1/34 = 0,532352941

90.4/240 = 0,376666667.

The limiting reagent is CuO.